OpenStudy (anonymous):

1.Find the compound interest on Rs.25000 FOR 6 months at 8%p.a when interest is compounded quaterly @Mathematics

OpenStudy (anonymous):

how many times interest will be paid to u ??

OpenStudy (anonymous):

rate of interest is 8 % p. a. so quaterly payment will be there

OpenStudy (anonymous):

what is the formula of compounding

OpenStudy (anonymous):

i know all these but still im asking u to help u it in better way

OpenStudy (anonymous):

A = P(1 + R/100 )^ t

OpenStudy (anonymous):

gr8 gary

OpenStudy (anonymous):

so R = 8 % P.A. but quaterly payment will be there right so what will be R then

OpenStudy (anonymous):

8/4 % P.A ?????????????

OpenStudy (anonymous):

gud

OpenStudy (anonymous):

but not 2 % P.A. it will be 2% per quater

OpenStudy (anonymous):

so ur formula can be re-written as \[Amount = Principal \times (1 + \frac{Rate}{4 \times 100})^t\]

OpenStudy (anonymous):

got it ???

OpenStudy (anonymous):

THNKS

OpenStudy (anonymous):

see the normal formula is \[Amount = Principal \times (1 + \frac{Rate}{n})^{nt}\]

OpenStudy (anonymous):

but in ur stated question their is only 2 times payment so we cant use "n"

OpenStudy (anonymous):

what happened ???

OpenStudy (anonymous):

\[Amount = 25000 \times (1 + 0.02)^{2}\]